The straight lines 3x(g+1)+4y(g-1)-3(g-1)=0 where g is a parameter, always pass through a fixed point (@,$) the value o of @+$ is

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
We can write it as 
 
g (3x + 4y - 3) + 1( 3x - 4y +3) = 0
 
Now solve these two equations which are in the form of L1 + k L2 = 0
 
x = 0 and y = ¾
 
Hence
 
@ + $ = 3/4